# Weyl law for 1-cycles

### v1

@@ Abstract.
We prove the Weyl law for the volume spectrum for $1$-cycles in $n$-dimensional manifolds which was conjectured by Gromov. We follow the strategy of Guth and Liokumovich of obtaining the Weyl law from parametric versions of the coarea inequality and the isoperimetric inequality. A version of the later for families of 00-cycles is shown in this article. We also obtain approximation results by $\delta$-localized families, which are used to prove the parametric inequalities.
@@ 1. Introduction
Let $(M,g)$ be an $n$-dimensional compact Riemannian manifold, possibly with boundary. In the 1960’s Almgren initiated a program of developing a Morse theory on the space $\mathcal{Z}_{k}(M;G)$ of flat cycles with coefficients in an abelian group $G$ (see [1],[2]). It follows from Almgren’s work that for each $10\leq k\leq n-$ 1 there is a non-trivial cohomology class $\lambda_{k}\in H^{n-k}(\mathcal{Z}_{k}(M;\mathbb{Z}_{2});\mathbb{Z}_{2})$ such that the following is true: a continuous $(n-k)$-parameter family of $k$-cycles $F:X\to\mathcal{Z}_{k}(M;\mathbb{Z}_{2})$ is a sweepout of $M$ if and only if $F^{*}(\lambda_{k})\neq 0$ in $H^{n-k}(X;\mathbb{Z}_{2})$ (see [11]). Moreover, all cup powers $\lambda_{k}^{p}$ are also non-trivial and we call the corresponding families of cycles $p$-sweepouts (namely, $F:X\to\mathcal{Z}_{k}(M;\mathbb{Z}_{2})$ is a $p$-sweepout if and only if $F^{*}(\lambda_{k}^{p})\neq 0$). We denote $\mathcal{P}_{p}^{k}$ the set of all $p$-sweepouts of $M$ by families of $k$-cycles. Corresponding to each $p$ we can define the $p$-width
@@
The $p$-widths correspond to volumes of certain generalized minimal submanifolds that arise via a min-max construction. In dimension $k=n-1$, $3\leq n\leq 7$, they are smooth minimal hypersurfaces [21],[22] and for $k=1$ they are stationary geodesic nets [21]. The study of $p$-widths led to many remarkable breakthroughs in recent years, including the resolution of Yau’s conjecture on existence of infinitely many minimal surfaces by Irie-Marques-Neves [13] in the generic case (see also [16] and [14] for similar results on stationary geodesic nets) and by Song [23] for arbitrary metrics. The related Multiplicity One Conjecture was proved by Chodosh-Mantoulidis [4] in the $3$-dimensional case and by Xin Zhou [25] and Marques-Neves [20] for $3\leq n\leq 7$. Gromov suggested to view these widths, or “volume spectrum”, as non-linear analogs of the eigenvalues of the Laplacian (see [7], [8], [9]). This framework is very useful for obtaining new results about widths and some other non-linear min-max geometric quantities. In line with this analogy Gromov conjectured that $\omega_{p}^{k}$’s satisfy an asymptotic Weyl law:
@@
for a certain universal constant $\alpha(n,k)>0$ depending only on $n$ and $k$. In [15], the conjecture was resolved for domains of $\mathbb{R}^{n}$ for arbitrary $n$ and $k$, and for $k=n-1$ for smooth compact manifolds. Later Guth and Liokumovich solved the case $k=1$, $n=3$ on compact manifolds in their article [12]. In the present work we solve the conjecture for $k=1$ and $n$ arbitrary.
@@
In [19], Marques, Neves and Song used the Weyl law for codimension $1$ cycles proved in [15] to show generic equidistribution of minimal hypersurfaces in ambient manifolds of dimension $3\leq n\leq 7$. Later, in a work with Xinze Li ([14]) we extended that result to the case $n=2$, proving generic equidistribution of immersed closed geodesics on surfaces. For that, we applied the following result of Chodosh and Mantoulidis which is proved in [4]: on closed surfaces, the $p$-widths $\omega_{p}^{1}(M,g)$ are realized by unions of immersed closed geodesics. When $n\geq 3$, they are realized by stationary geodesic nets, which are embedded graphs in $M$ that are stationary with respect to the length functional. One could ask whether an equidistribution result like the one showed by Marques, Neves and Song holds for $k$-dimensional minimal surfaces (possibly with singularities) in $n$-dimensional manifolds, $k<n-1$. In [14], we solved the case $k=1$ and $n=3$ using the Weyl law for $1$-cycles in $3$-manifolds from [12]; proving generic equidistribution of stationary geodesic nets in $3$-manifolds. The only reason for the restriction to $n=3$ was that the Weyl law for $1$-cycles was still open for $n\geq 4$. Using Theorem 1.1 and [14][Theorem 1.5], we can now extend that result to $n\geq 4$, obtaining the following.
@@
The proof of Theorem 1.1 is discussed in Section 5. We follow the strategy proposed in [12] which consists of obtaining the Weyl law from parametric versions of the Isoperimetric Inequality and of the Coarea Inequality. For that purpose, we prove the following theorem, which solves [12][Conjecture 1.3] for $k=0$.
Theorem 1.3 (Parametric Isoperimetric Inequality for 00-cycles).
Let $(M,g)$ be a compact $n$-dimensional Riemannian manifold, possibly with boundary. Then there exists a constant $C=C(M,g)$ such that the following holds. Let let $X=X^{p}$ be a cubical complex of dimension $p$ and let $F:X^{p}\to\mathcal{Z}_{0}(M;\mathbb{Z}_{2})$ be a contractible family of 00-cycles, $\dim(X)=p$. Denote
@@
The proof of Theorem 1.3 is discussed in Section 4. In Section 4.1, we prove the result for families on the Euclidean disk $\mathbb{D}^{n}$ and then in Section 4.3 we extend it to any manifold $M$ by considering a fine triangulation $\{Q_{i}\}_{1\leq i\leq N}$ of $M$ and applying the result for the disk to each $F_{i}(x)=F(x)\llcorner Q_{i}$. The previous reduces the problem to filling a family in the $(n-1)$-skeleton of $M$, allowing us to argue by induction. Breaking the family $F$ into the pieces $F_{i}(x)$ yields the following obstacle: $\operatorname{\mathbf{M}}(F_{i}(x)\llcorner\partial Q_{i})$ might be very large, making the fillings on each $Q_{i}$ very long. In Section 4.2, we prove Proposition 4.10 which states that each $F_{i}$ can be slightly perturbed to a continuous family of absolute 00-cycles $F_{i}^{\prime}$ such that $\operatorname{\mathbf{M}}(F_{i}^{\prime}(x)\llcorner B)$ is controlled for some small ball $B\subseteq\partial Q_{i}$, which turns out to be enough to resolve the previously mentioned obstacle.
To prove the Weyl law, we also use the Parametric Coarea Inequality for $1$-cycles in $n$-manifolds, which is proved in [24]. It solves [12][Conjecture 1.6] in the case $k=1$ and $n$ arbitrary. We state it below.
@@
Let $(M,g)$ be a compact $n$-dimensional Riemannian manifold with boundary, $n\geq 4$. Let $\alpha\in(0,1)$. There exists $p_{0}=p_{0}(M,g,\alpha)$ and a constant $C=C(\partial M)$ such that the following is true. Let $p\geq p_{0}$, let $X=X^{p}$ be a cubical complex of dimension $p$ and let $F:X^{p}\to\mathcal{Z}_{1}(M,\partial M)$ be a family of relative $1$-cycles which is continuous in the flat topology and has no concentration of mass. Then for every $\varepsilon>0$, there exists a continuous family $F^{\prime}:X^{p}\to\mathcal{I}_{1}(M)$ of absolute chains in $M$ such that
@@
    $\operatorname{spt}(\partial F^{\prime}(x))\subseteq\partial M$. Therefore $F^{\prime}$ induces a family of relative $1$-cycles $F^{\prime}:X^{p}\to\mathcal{Z}_{1}(M,\partial M)$ which is denoted in the same way.
@@
    $\mathcal{F}(F(x),F^{\prime}(x))\leq\varepsilon$ as relative cycles in the space $\mathcal{Z}_{1}(M,\partial M)$.
@@
    $\operatorname{\mathbf{M}}(F^{\prime}(x))\leq(1+\gamma_{p})\operatorname{\mathbf{M}}(F(x))+C(\partial M)p^{\frac{n-2}{n-1}+\alpha}$.
@@
    $\operatorname{\mathbf{M}}(\partial F^{\prime}(x))\leq C(\partial M)p^{1+\alpha}$.
In the previous $(\gamma_{p})_{p\in\mathbb{N}}$ is a sequence of positive real numbers converging to 00 as $p\to\infty$.
In order to obtain the parametric inequalities, it is convenient to work with $\delta$-localized families, which were first defined in [12][Section 2]. These are families $F:X\to\mathcal{Z}_{k}(M,\partial M;G)$ such that for every cell $C$ of $X$ and for every $x,y\in C$, $F(x)-F(y)$ is supported in a collection of balls with sum of radii less than $\delta$ (see Definition 3.2). $G$ can be taken to be $\mathbb{Z}$ or $\mathbb{Z}_{p}$ for some prime $p$. Having this control over how the family $F$ behaves on the cells of $X$ allows us to produce new families with desired mass bounds and additional properties. For example, it permits us to construct families inductively skeleton by skeleton, starting from a discrete family defined in the 00-skeleton of $X$, keeping track of quantities like the masses of the chains involved, of its boundaries, $\mathcal{F}(F(x)-F(y))$ for $x,y\in C$ and the supports of the $F(x)$. In order to use these families for our constructions, we need to show that every continuous family can be arbitrarily well approximated by a $\delta$-localized one. This is proved both for families of cycles and for families of chains, as stated below.
@@
Let $G=\mathbb{Z}$ or $\mathbb{Z}_{p}$ and let $F:X\to\mathcal{Z}_{k}(M,\partial M;G)$ be a continuous map without concentration of mass. Let $\varepsilon,\delta>0$. Then there exists a continuous map $F^{\prime}:X\to\mathcal{Z}_{k}(M,\partial M;G)$ without concentration of mass and a fine cubulation of $X$ such that
@@
Let $G=\mathbb{Z}$ or $\mathbb{Z}_{p}$ and let $G:X\to\mathcal{I}_{k+1}(M,\partial M;G)$ be a continuous map without concentration of mass. Denote $F=\partial G$ and let $\varepsilon,\delta>0$. Then there exists a continuous map $G^{\prime}:X\to\mathcal{I}_{k+1}(M,\partial M;G)$ without concentration of mass and a fine cubulation of $X$ such that
@@
In order to prove the previous theorems, we first obtain versions of them for discrete families $F:X_{0}\to\mathcal{Z}_{k}(M,\partial M;G)$ (resp. $\mathcal{I}_{k+1}(M,\partial M;G)$) and then use the techniques to go from discrete to continuous families developed in [12][Section 2]. All this is discussed in Section 3.
Acknowledgments. I am very grateful to Yevgeny Liokumovich for suggesting this problem and for all the valuable conversations we have had while I was working on it.
@@ 2. Preliminaries
2.1. Triangulated Riemannian manifolds with corners
We will need to work with a class of Riemannian manifolds with piecewise smooth boundary which admit a compatible PL structure. For that purpose, we introduce the following definitions, which are inspired in [12][Section 2.1] and [3].
@@ 2.1. Triangulated Riemannian manifolds with corners
Let $\Theta$ be a simplicial complex. An Euclidean structure on $\Theta$ is the path metric obtained by giving each simplex the structure of an Euclidean simplex. In other words, each $j$-simplex $\Delta\in\Theta$ is isometric to a linear $j$-simplex embedded in $\mathbb{R}^{n}$ (not necessarily the standard one) and the metric $d$ in $\Theta$ is the corresponding path metric (which comes from infimizing lengths of paths connecting two given points which are piecewise supported in different simplices $\Delta\in\Theta$).
@@
Let $\mu\geq 1$. A $\mu$-triangulated Riemannian manifold with corners $(M,g)$ is a compact manifold with corners $M$ provided with a triangulation $\tau:\Theta\to M$ and a smooth Riemannian metric $g$ such that $\Theta$ is an Euclidean simplicial complex and $\tau:\Theta\to M$ is $\mu$-bilipschitz. We say $M$ is a triangulated Riemannian manifold with corners if it is a $\mu$-triangulated Riemannian manifold with corners for some $\mu\geq 1$. A triangulation $\tau:\Theta\to M$ with the previous properties is called a $\mu$-triangulation.
@@
By [3][Theorem 1.2 and 1.3], every compact Riemannian manifold (possibly with boundary) can be $\mu$-triangulated for every $\mu>1$. Therefore along the paper we will consider $\mu$-triangulated Riemannian manifolds with corners with $\mu$ very close to $1$.
@@
Let $(M,g)$ be a $\mu$-triangulated Riemannian manifold with corners. We denote $\Sigma=\partial M$ and $d$ the metric on $M$ induced by the Riemannian metric $g$. Given two subsets $A,B\subseteq M$ we define
@@
Let $(M,g)$ be a $\mu$-triangulated Riemannian manifold with corners and let $A\subseteq M$ be a subcomplex of a certain refinement of the triangulation $\tau:\Theta\to M$ (in particular, $A$ is a PL submanifold of $M$). For each $r>0$, we denote
$N_{r}A=\{x\in M:\operatorname{dist}(x,A)<r\}.$
Definition 2.6.
Given a $\mu$-triangulated Riemannian manifold with corners $(M,g)$, we consider $r_{0}=r_{0}(M,g,\mu)$ and a $\mu$-bilpschitz map $\operatorname{Exp}:\Sigma\times[0,r_{0}]\to M$ such that
  1. (1)
    $\operatorname{Exp}(x,0)=x$ for every $x\in\Sigma$ and
  2. (2)
    $N_{\mu^{-1}r_{0}}\Sigma\subseteq\operatorname{Exp}(\Sigma\times[0,r_{0}])\subseteq N_{\mu r_{0}}\Sigma$
as constructed in [12][Section 2.1]. This maps plays the role of the exponential map in a Riemannian manifold with smooth boundary.

### v2

@@ Abstract.
We prove the Weyl law for the volume spectrum for $1$-cycles in $n$-dimensional manifolds which was conjectured by Gromov. We follow the strategy of Guth and Liokumovich of obtaining the Weyl law from parametric versions of the coarea inequality and the isoperimetric inequality. A version of the later for families of 0-cycles is shown in this article. We also obtain approximation results by $\delta$-localized families, which are used to prove the parametric inequalities.
@@ 1. Introduction
Let $(M,g)$ be an $n$-dimensional compact Riemannian manifold, possibly with boundary. In the 1960’s Almgren initiated a program of developing a Morse theory on the space $\mathcal{Z}_{k}(M;G)$ of flat cycles with coefficients in an abelian group $G$ (see [FA62],[FA65]). It follows from Almgren’s work that for each $0\leq k\leq n-1$ there is a non-trivial cohomology class $\lambda_{k}\in H^{n-k}(\mathcal{Z}_{k}(M;\mathbb{Z}_{2});\mathbb{Z}_{2})$ such that the following is true: a continuous (nk)(n-k)-parameter family of $k$-cycles $F:X\to\mathcal{Z}_{k}(M;\mathbb{Z}_{2})$ is a sweepout of $M$ if and only if $F^{*}(\lambda_{k})\neq 0$ in $H^{n-k}(X;\mathbb{Z}_{2})$ (see [GuthMinMax]). Moreover, all cup powers $\lambda_{k}^{p}$ are also non-trivial and we call the corresponding families of cycles $p$-sweepouts (namely, $F:X\to\mathcal{Z}_{k}(M;\mathbb{Z}_{2})$ is a $p$-sweepout if and only if $F^{*}(\lambda_{k}^{p})\neq 0$). We denote $\mathcal{P}_{p}^{k}$ the set of all $p$-sweepouts of $M$ by families of $k$-cycles. Corresponding to each $p$ we can define the $p$-width
@@
The $p$-widths correspond to volumes of certain generalized minimal submanifolds that arise via a min-max construction. In dimension k=n1k=n-1, $3\leq n\leq 7$, they are smooth minimal hypersurfaces [JP81],[SchoenSimon] and for $k=1$ they are stationary geodesic nets [JP73]. The study of $p$-widths led to many remarkable breakthroughs in recent years, including the resolution of Yau’s conjecture on existence of infinitely many minimal surfaces by Irie-Marques-Neves [IMN] in the generic case (see also [LS] and [LiSta] for similar results on stationary geodesic nets) and by Song [Song] for arbitrary metrics. The related Multiplicity One Conjecture was proved by Chodosh-Mantoulidis [ChoMan20] in the $3$-dimensional case and by Xin Zhou [ZhouMult1] and Marques-Neves [MNMult1] for $3\leq n\leq 7$. Gromov suggested to view these widths, or “volume spectrum”, as non-linear analogs of the eigenvalues of the Laplacian (see [Gromov86], [Gromov02], [Gromov09]). This framework is very useful for obtaining new results about widths and some other non-linear min-max geometric quantities. In line with this analogy Gromov conjectured that $\omega_{p}^{k}$’s satisfy an asymptotic Weyl law:
@@
for a certain universal constant $\alpha(n,k)>0$ depending only on $n$ and $k$. In [LMN], the conjecture was resolved for domains of $\mathbb{R}^{n}$ for arbitrary $n$ and $k$, and for k=n1k=n-1 for smooth compact manifolds. Later Guth and Liokumovich solved the case $k=1$, $n=3$ on compact manifolds in their article [GL22]. In the present work we solve the conjecture for $k=1$ and $n$ arbitrary.
@@
In [MNS], Marques, Neves and Song used the Weyl law for codimension-$1$ cycles proved in [LMN] to show generic equidistribution of minimal hypersurfaces in ambient manifolds of dimension $3\leq n\leq 7$. Later, in a work with Xinze Li ([LiSta]) we extended that result to the case $n=2$, proving generic equidistribution of immersed closed geodesics on surfaces. For that, we applied the following result of Chodosh and Mantoulidis which is proved in [ChoMan]: on closed surfaces, the $p$-widths $\omega_{p}^{1}(M,g)$ are realized by unions of immersed closed geodesics. When $n\geq 3$, they are realized by stationary geodesic nets, which are embedded graphs in $M$ that are stationary with respect to the length functional. One could ask whether an equidistribution result like the one showed by Marques, Neves and Song holds for $k$-dimensional minimal surfaces (possibly with singularities) in $n$-dimensional manifolds, k<n1k<n-1. In [LiSta], we solved the case $k=1$ and $n=3$ using the Weyl law for $1$-cycles in $3$-manifolds from [GL22]; proving generic equidistribution of stationary geodesic nets in $3$-manifolds. The only reason to restrict ourselves to $n=3$ was that the Weyl law for $1$-cycles was still open for $n\geq 4$. Using Theorem 1.1 and [LiSta][Theorem 1.5], we can now extend that result to $n\geq 4$, obtaining the following.
@@
The proof of Theorem 1.1 is discussed in Section 5. We follow the strategy proposed in [GL22] which consists of obtaining the Weyl law from parametric versions of the Isoperimetric Inequality and of the Coarea Inequality. For that purpose, we prove the following theorem, which solves [GL22][Conjecture 1.3] for $k=0$.
Theorem 1.3 (Parametric Isoperimetric Inequality for 0-cycles).
Let $(M,g)$ be a compact $n$-dimensional Riemannian manifold (possibly with boundary), $n\geq 2$. Then there exists a constant $C=C(M,g)$ such that the following holds. Let $X=X^{p}$ be a cubical complex of dimension $p$ and let $F:X^{p}\to\mathcal{Z}_{0}(M;\mathbb{Z}_{2})$ be a contractible family of 0-cycles, $\dim(X)=p$. Denote
@@
The proof of Theorem 1.3 is discussed in Section 4. In Section 4.1, we prove the result for families on the Euclidean disk $\mathbb{D}^{n}$ and then in Section 4.3 we extend it to any manifold $M$ by considering a fine triangulation $\{Q_{i}\}_{1\leq i\leq N}$ of $M$ and applying the result for the disk to families $F_{i}(x)$ with $\operatorname{spt}(F_{i}(x))\subseteq Q_{i}$ and $\sum_{i=1}^{N}F_{i}(x)=F(x)$. The previous reduces the problem to filling a family in the (n1)(n-1)-skeleton of $M$, allowing us to argue by induction. Breaking the family $F$ into continuous pieces $F_{i}(x)$ supported in the $Q_{i}$ yields the following obstacle: $\operatorname{\mathbf{M}}(F_{i}(x)\llcorner\partial Q_{i})$ might be very large, making the fillings on each $Q_{i}$ very long. In Section 4.2, we prove Proposition 4.10 which states that each $F_{i}$ can be slightly perturbed to a continuous family of absolute 0-cycles $F_{i}^{\prime}$ such that $\operatorname{\mathbf{M}}(F_{i}^{\prime}(x)\llcorner B)$ is controlled for some small ball $B\subseteq\partial Q_{i}$, which turns out to be enough to resolve the previously mentioned obstacle.
To prove the Weyl law, we also use the Parametric Coarea Inequality for $1$-cycles in $n$-manifolds, which is proved in [StaCoarea]. It resolves [GL22][Conjecture 1.6] in the case $k=1$ and $n$ arbitrary. We state it below.
@@
Let $M$ be a compact $n$-dimensional Riemannian manifold with boundary, $n\geq 4$. Let $\alpha\in(0,1)$. There exist $p_{0}=p_{0}(M,\alpha)$, a constant $c=c(n)$ and a sequence of numbers $\gamma_{p}=\gamma_{p}(n,\alpha)$ converging to 0 such that the following is true. Let $p\geq p_{0}$ and let $F:X^{p}\to\mathcal{Z}_{1}(M,\partial M;\mathbb{Z}_{2})$ be a family of relative $1$-cycles which is continuous in the flat topology and has no concentration of mass. Then for every $\varepsilon>0$, there exists a continuous family $F^{\prime}:X^{p}\to\mathcal{I}_{1}(M;\mathbb{Z}_{2})$ of absolute chains in $M$ such that
@@
    $\operatorname{spt}(\partial F^{\prime}(x))\subseteq\partial M$. Therefore $F^{\prime}$ induces a family of relative $1$-cycles $F^{\prime}:X^{p}\to\mathcal{Z}_{1}(M,\partial M;\mathbb{Z}_{2})$ which is denoted in the same way.
@@
    $\mathcal{F}(F(x),F^{\prime}(x))\leq\varepsilon$ as relative cycles in the space $\mathcal{Z}_{1}(M,\partial M;\mathbb{Z}_{2})$.
@@
    $\operatorname{\mathbf{M}}(F^{\prime}(x))\leq(1+\gamma_{p})\operatorname{\mathbf{M}}(F(x))+c\operatorname{Vol}_{n-1}(\partial M)p^{\frac{n-2}{n-1}+\alpha}$.
@@
    $\operatorname{\mathbf{M}}(\partial F^{\prime}(x))\leq c\operatorname{Vol}_{n-1}(\partial M)p^{1+\alpha}$.
In order to obtain the parametric inequalities, it is convenient to work with $\delta$-localized families, which were first defined in [GL22][Section 2]. These are families $F:X\to\mathcal{Z}_{k}(M,\partial M;\mathbb{G})$ such that for every cell $C$ of $X$ and for every $x,y\in C$, F(x)F(y)F(x)-F(y) is supported in a collection of balls with sum of radii less than $\delta$ (see Definition 3.3). $\mathbb{G}$ can be taken to be $\mathbb{Z}$ or $\mathbb{Z}_{p}$ for some prime $p$. Having this control over how the family $F$ behaves on the cells of $X$ allows us to produce new families with desired mass bounds and additional properties. For example, it permits us to construct families inductively skeleton by skeleton, starting from a discrete family defined in the 0-skeleton of $X$, keeping track of quantities like the masses of the chains involved, of its boundaries, $\mathcal{F}(F(x)-F(y))$ for $x,y\in C$ and the supports of the $F(x)$. In order to use these families for our constructions, we need to show that every continuous family can be arbitrarily well approximated by a $\delta$-localized one. This is proved both for families of cycles and for families of chains, as stated below.
@@
Let $\mathbb{G}=\mathbb{Z}$ or $\mathbb{Z}_{p}$ and let $F:X\to\mathcal{Z}_{k}(M,\partial M;\mathbb{G})$ be a continuous map without concentration of mass. Let $\varepsilon,\delta>0$. Then there exists a continuous map $F^{\prime}:X\to\mathcal{Z}_{k}(M,\partial M;\mathbb{G})$ without concentration of mass and a fine cubulation of $X$ such that
@@
Let $\mathbb{G}=\mathbb{Z}$ or $\mathbb{Z}_{p}$ and let $G:X\to\mathcal{I}_{k+1}(M,\partial M;\mathbb{G})$ be a continuous map without concentration of mass. Denote $F=\partial G$ and let $\varepsilon,\delta>0$. Then there exists a continuous map $G^{\prime}:X\to\mathcal{I}_{k+1}(M,\partial M;\mathbb{G})$ without concentration of mass and a fine cubulation of $X$ such that if $F^{\prime}=\partial G^{\prime}$ then
@@
In order to prove the previous theorems, we first obtain versions of them for discrete families $F:X_{0}\to\mathcal{Z}_{k}(M,\partial M;\mathbb{G})$ (resp. $\mathcal{I}_{k+1}(M,\partial M;\mathbb{G})$) and then use the techniques to go from discrete to continuous families developed in [GL22][Section 2]. All this is discussed in Section 3.
Acknowledgments. I am very grateful to Yevgeny Liokumovich for suggesting this problem and for all the valuable conversations we have had while I was working on it. I would also like to express my deep gratitude to the Hausdorff Research Institute for Mathematics, where part of this work was completed during my participation in the Trimester Program: Metric Analysis, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.
@@ 2. Preliminaries
2.1. Piecewise smooth Riemannian manifolds
We will need to work with a class of Riemannian manifolds with piecewise smooth boundary which admit a compatible PL structure. For that purpose, we introduce the following definitions.
@@ 2.1. Piecewise smooth Riemannian manifolds
A Euclidean polyhedron $P$ is a linearly embedded finite simplicial complex in $\mathbb{R}^{N}$ for some $N\in\mathbb{N}$.
@@
Let $P$ be a Euclidean polyhedron. A map $\phi:P\to\mathbb{R}^{N}$ is piecewise smooth if $\phi$ is continuous and for each face $F$ of a certain triangulation $T$ of $P$ the map $\phi|_{F}$ is smooth. $\phi$ is a piecewise smooth embedding if it is a topological embedding and $\phi|_{F}$ is a smooth embedding for each face $F$ of $P$.
Definition 2.3.
An $n$-dimensional Riemannian polyhedron is a is a piecewise smoothly embedded $n$-dimensional polyhedron in $\mathbb{R}^{N}$ (i.e. the image of a Euclidean polyhedron under a piecewise smooth embedding into $\mathbb{R}^{n}$ for some value of $N$) equipped with the piecewise smooth Riemannian metric $g$ induced by the Euclidean metric in $\mathbb{R}^{N}$.
Definition 2.4.
Let $P$ be a Riemannian polyhedron embedded in $\mathbb{R}^{N}$ and denote the corresponding embedding by $\iota:P\to\mathbb{R}^{N}$. A smooth triangulation of $P$ is a homeomorphism $\tau:P^{\prime}\to P$ where $P^{\prime}$ is a Euclidean polyhedron provided with a triangulation $T^{\prime}$ and $\iota\circ\tau:P^{\prime}\to\mathbb{R}^{n}$ is a piecewise smooth embedding with respect to the triangulation $T^{\prime}$. A triangulated Riemannian polyhedron is a pair $(P,T)$ where $P$ is a Riemannian polyhedron and $T$ is the triangulation on $P$ induced by the $T^{\prime}$ just defined. A refinement of $T$ is the triangulation induced by taking a refinement of the corresponding $T^{\prime}$.
@@
Riemannian polyhedra (and hence Euclidean ones) admit different triangulations. For example, we can obtain infinitely many of them from a single one by refinement.
@@
A piecewise smooth Riemannian manifold $(M^{n},g)$ is an $n$-dimensional Riemannian polyhedron provided with a topological manifold with boundary structure (i.e. with an atlas of class $C^{0}$). In other words, it is an $n$-dimensional topological submanifold with boundary of $\mathbb{R}^{N}$ (for some $N\in\mathbb{N}$) which admits a smooth triangulation and is equipped with a continuous Riemannian metric $g$ which is smooth on each simplex of that triangulation.
Definition 2.7.
A piecewise smooth Riemannian manifold $(M,g)$ is almost $1$-Lipchitz triangulable if for every $\varepsilon>0$ there exists a smooth $(1+\varepsilon)$-bilipschitz triangulation $\tau_{\varepsilon}:P_{\varepsilon}\to M$ (here $P_{\varepsilon}$ is a Euclidean polyhedron according to Definition 2.4).
Remark 2.8.
By the Nash Embedding Theorem and Theorems 1.1 and 1.2 of [Bowditch], every smooth compact Riemannian manifold with boundary is almost $1$-Lipschitz triangulable. So is every piecewise linear submanifold of $\mathbb{R}^{N}$. Thus the collection of all almost $1$-Lipschitz triangulable piecewise smooth Riemannian manifolds includes both the category of compact smooth manifolds with boundary and the category of compact PL manifolds with boundary. It will be the correct set up for our constructions because we are interested in proving results about compact Riemannian manifolds but our methods will involve considering $(1+\varepsilon)$-bilipschitz triangulations.
Remark 2.9.
Because of the works of Akopyan [Akopyan] and Minemyer [Minemyer], every Euclidean simplicial complex as defined by Bowditch in [Bowditch] can be piecewise linearly embedded in $\mathbb{R}^{N}$ and hence can be regarded as a Euclidean polyhedron.
Remark 2.10.
As far as the author knows, the question whether every piecewise smooth Riemannian manifold is almost $1$-Lipschitz triangulable has not been studied. Maybe the methods of Bowditch [Bowditch] also work in this case.
Remark 2.11.
Notice that in Definition 2.7, $P_{\varepsilon}$ inherits a topological manifold structure from $M$ and hence it is a compact PL manifold with boundary. So the condition of being almost $1$-Lipschitz triangulable can be re-expressed as being $(1+\varepsilon)$-bilipschitz diffeomorphic to a compact PL manifold with boundary for every $\varepsilon>0$.
Definition 2.12.
Let $(M,g)$ be a piecewise smooth Riemannian manifold. We denote $\Sigma=\partial M$ (which is a subpolyhedron of $M$) and $d$ the metric on $M$ induced by the Riemannian metric $g$. Given two subsets $A,B\subseteq M$ we define
@@
Let $A\subseteq M$ be a subset of $M$. For each $r>0$, we denote
$N_{r}A$ $=\{x\in M:\operatorname{dist}(x,A)<r\}$
$D_{r}A$ $=\{x\in M:\operatorname{dist}(x,A)=r\}.$
@@
Given a smooth Riemannian manifold with boundary $(M,g)$, we denote its exponential map by $\operatorname{Exp}$.